If consumption function is given as c= $25+0.8Y where yd=y-t ,tax rate=10% investment=$50 calculate equilibrium level of income as well as the government expenditure multiplier.

1
Expert's answer
2021-12-10T11:25:08-0500

At equilibrium, Y=C+I. Therefore, we have


"Y=25+0.8(Y-0.1Y)+50\\[0.3cm]
Y=75+0.64Y\\[0.3cm]
Y-0.64Y=75\\[0.3cm]
0.36Y=75\\[0.3cm]
Y^*=\boxed{\$208.33}"

The government expenditure multiplier is equal to


"\dfrac{\Delta Y}{\Delta G}=\dfrac{1}{1-mpc}\\[0.3cm]
\dfrac{\Delta Y}{\Delta G}=\dfrac{1}{1-0.8}\\[0.3cm]
\dfrac{\Delta Y}{\Delta G}=\boxed{5}"